90% acid is to be added to a 15% solution to get 18 liters of a 25% solution. What amounts of each should be used?
Answer by longjonsilver(2297)

Let x be the amount of the 90% solution.
So how much of the 15% solution is there? Answer is 18-x

and BEFORE = AFTER

So,
BEFORE: 0.90x + 0.15(18-x)
AFTER: 0.25*18

--> 0.90x + 0.15(18-x) = 0.25*18
0.90x + 2.7 - 0.15x = 4.5
0.75x + 2.7 = 4.5
0.75x = 1.8
x = 2.4 litres

so, 2.4 litres of the 90% solution and 15.6 litres of the 15% solution

jon.